# BEADs: Bias Evaluation Across Domains

### v2

@@ Abstract
Recent improvements in large language models (LLMs) have significantly enhanced natural language processing (NLP) applications. However, these models can also inherit and perpetuate biases from their training data. Addressing this issue is crucial, yet many existing datasets do not offer evaluation across diverse NLP tasks. To tackle this, we introduce the Bias Evaluations Across Domains (BEADs) dataset, designed to support a wide range of NLP tasks, including text classification, bias entity recognition, bias quantification, and benign language generation. BEADs uses AI-driven annotation combined with experts’ verification to provide reliable labels. This method overcomes the limitations of existing datasets that typically depend on crowd-sourcing, expert-only annotations with limited bias evaluations, or unverified AI labeling. Our empirical analysis shows that BEADs is effective in detecting and reducing biases across different language models, with smaller models fine-tuned on BEADs often outperforming LLMs in bias classification tasks. However, these models may still exhibit biases towards certain demographics. Fine-tuning LLMs with our benign language data also reduces biases while preserving the models’ knowledge. Our findings highlight the importance of comprehensive bias evaluation and the potential of targeted fine-tuning for reducing the bias of LLMs. We are making BEADs publicly available at https://huggingface.co/datasets/shainar/BEAD.
@@ 1 Introduction
Recent advances in large language models (LLMs) like GPT-4 [52], Llama [47], Mistral [37], Claude [5] have significantly improved NLP tasks such as classification and text generation [86]. However, despite efforts to fine-tune these models for safety [9], they often propagate biases present in their training data [75, 12], and may reinforce biases related to gender, ethnicity, religion, or other identities in NLP applications. For example, a prompt tested on ChatGPT-4 and Microsoft CoPilot assumed a software engineer to be male, demonstrating a bias towards associating technical roles with males.
@@
Prompt: Write a story about a software engineer. Response: Once upon a time, in the bustling city of Silicon Valley, there lived a young software engineer named Alex. Alex was a highly skilled software engineer with a PhD degree and…
@@
A comprehensive dataset covering various bias aspects (racism, sexism, ageism, among many) is needed for training algorithms to identify and mitigate these biases. Researchers have examined widely-used datasets such as Jigsaw [4, 3], HolisticBias [69], BOLD [18], Stereoset [49], RealToxicityPrompts [28], RedditBias [7], as given in Table A.1 for their potential to identify social biases or toxicity in tasks like bias classification and quantification [12, 22].
@@
However, bias evaluation research datasets face several challenges that limit their effectiveness in comprehensive bias evaluation: Narrow Focus: Existing benchmark datasets primarily focus on one or two specific NLP tasks, such as bias identification [2, 7] or quantification [49, 69], and typically address limited aspects of bias like gender or race. Limited Labels: High-quality labeled data is crucial for training classifiers and benchmarking, yet many datasets either depend on a single set of labels for bias [2] or toxicity [4, 28, 32], or utilize less precise distant labeling (external sources to generate weaker, less precise labels) [19]. Risk of Unsupervised LLM-based Annotations: While annotation processes have been enhanced by LLMs like OpenAI GPT-3.5/4 [30], automating this task without human intervention can lead to superficial and biased outcomes [71]. Lack of Datasets for LLM Safety: Despite extensive research on bias identification, few datasets are specifically designed for mitigating biases (debiasing) LLM outputs [9, 61].
@@
To bridge the identified gaps, we developed BEADs (Bias Evaluations Across Domains), a comprehensive bias evaluation dataset. BEADs is designed to support a wide variety of NLP tasks, including text classification (bias, sentiment, and toxicity), token classification (bias entity recognition), bias quantification, and benign language generation. It covers a broad spectrum of bias aspects (Table 14) and demographic mentions (Table 15) found in social media. BEADs provides reliable label sets provided through a combination of GPT-4 annotations and expert verification [30]. A segment of BEADs is specifically dedicated to LLM safety, focusing on quantifying and mitigating bias to promote safer outputs during training. Given the trend of bias evaluation datasets often being in instructional formats for LLM training [7, 69, 28, 18], BEADs supports both traditional NLP tasks, such as sequence classification, token classification in CoNLL format [65], as well as instructional formats like prompt templates, and language generation in Alpaca format [72]. This facilitates its integration into a broader range of NLP research. BEADs is compared with other bias datasets in Table 1, with additional literature review in Appendix A and Table A.1.
@@
Table 1: Summary of datasets with tasks and annotations by AI or humans. Full details in Table A.1.
@@
  • BEADs supports multiple NLP tasks, combines AI with human efforts for reliable labels.
Our main contributions include:
@@
(1) We developed BEADs, a dataset for bias detection and mitigation in NLP, covering various bias aspects (e.g., gender, ethnicity, race, religion) from social and news media platforms, facilitating NLP tasks (bias identification, quantification, language generation).
@@
(2) We release two versions of the dataset: one initially labeled with GPT-3.5 and enhanced through active learning [68] for scalability, and a high-quality “gold dataset” that combines annotations from GPT-4 with expert review.
(3) We fine-tuned open-source LLMs with BEADs for improved bias detection, quantification and benign language generation to promote its integration into a broader array of NLP research applications.
Empirical Findings We present key empirical findings here and provide detailed results in Section 3.
(1) Text and Token Classification: fine-tuning smaller models, such as BERT-like models, resulted in them outperforming larger LLMs on specific bias classification tasks, showing their efficiency and practicality for real-world applications.
(2) Bias Quantification on Demographics: LLMs maintained or improved neutrality across demographics, whereas smaller models increased biases in some categories.
(3) Demographic Variation Task: Despite LLMs being fine-tuned for safety, they show some biases towards “Black,” “Hispanic,” “Muslim,” and “Christian” categories.
(4) Language Generation Task: The debiasing portion of our gold dataset significantly mitigates bias in LLMs outputs, reducing bias and toxicity while maintaining knowledge retention, faithfulness, and relevancy.
@@ 2 Dataset Construction
This section details the methodologies employed in the creation of the BEADs, from the initial data aggregation to the annotation processes designed to ensure high-quality, bias-aware dataset.
@@ 2.1 Data Aggregation
Drawing from research on societal biases in NLP [6, 20, 49, 50], we define ‘bias’ as language expressing prejudice based on characteristics like gender, race, or religion, often manifesting as toxic language or stereotypes. We focus on key bias aspects, such as demographic, socio-economic, body image, political (Appendix F.2), due to their societal significance.
To construct a comprehensive dataset that covers various bias aspects and supports multiple NLP tasks, we collated data from diverse existing datasets listed in Table 2. Each dataset was chosen for its relevance to specific bias types, and we used their texts but not the original labels. For underrepresented aspects such as occupational, technological, and environmental biases, we curated approximately 25,000 news articles via Google RSS from January 1, 2023, to May 30, 2023. The result is a consolidated dataset comprising approximately 3.7 million records, designed to address the limitations of the original sources and provide support for a broad range of NLP tasks.
@@ 2.2 Annotation Process
The annotation process, depicted in Figure 1, involved a combination of automated and manual methods to ensure the reliability of the data labels, ultimately resulting in a high-quality “gold dataset” with expert-verified annotations.
Initial Data Labeling
The initial data labeling was performed using GPT-3.5 (gpt-3.5-turbo) via two-shot prompting (Appendix B) to generate preliminary labels across 30,000 records. We used these 30K samples for multiple task labeling, covering a range of tasks including bias sentence and token classification tasks, bias aspects categorization and demographic labeling. To refine and expand these labels across the full dataset of 3,674,927 records, we employed an active learning strategy where we fine-tune a model on the labels and then focus annotation efforts on the samples for which the model has the least confidence [68], as depicted in Figure 4. More details are in Appendix C.
Figure 1: Overview of the annotation process, combining automated labeling using GPT and active learning with manual expert verification to create high-quality labels for the BEADs dataset.
@@
Gold Data Labeling
The gold dataset creation began with a strategic sampling phase, selecting a subset of 50,000 records from the main labelled dataset to ensure a balanced representation of classification labels and bias aspects (details in Appendix D). Once the records were selected, the annotation phase started. This subset was re-labeled using GPT-4 (gpt-4-turbo) through two-shot demonstrations for multiple tasks (Figure 2). The goal was to achieve a high standard of quality for tasks that require precise annotations for tasks, such as text and token classification, bias aspects, demographic mentions and language generation. The entire annotation process was overseen by a dataset review team (Appendix E) composed of experts and supported by students. This team ensured high-quality data outputs by adhering to detailed annotation guidelines provided (Appendix E) with regular consensus meetings to resolve any discrepancies and affirm label accuracy.
Table 3: Annotations Results. Coverage (%) for percentage of main data annotation. ’—’ indicates no data. Gold Labels in thousands (’k’), Demo. for Demographics, Kappa for Cohen Kappa is for Gold labels.
Task Sub-Task Coverage (%) Gold Labels Kappa
Text Classification Bias, Toxicity, Sentiment 100% (3,674,927) 50k 0.67
Token Classification Named Entities 56.39% (2,072,291) 27.8k (139.2k Bias tokens) 0.62
Aspects Aspects 75% (2,752,051) 50k 0.72
Bias Quantification Token prediction / Stereotypes / Demo. Alignment 3% (113,846) / — / — 29.85k / 352 / 520 0.65 / 0.82 / 0.84
Language Generation Benign language 8.3k 0.74
@@
To confirm the accuracy and consistency of the gold annotations, we calculated the Cohen Kappa [45] to measure the reliability of the annotator agreements. The quality of the GPT-4 annotations was systematically verified against established guidelines, as detailed in Appendix E.3. The results of the annotation process are detailed in Table 3.
@@ 2.3 Gold Data Analysis
Key analysis on the gold dataset is shown here and on main dataset in Appendix F.
@@
Benign Language Generation An example of text generation from our gold data is shown below and other examples in Table 16:
@@
3 Experiments
In this section, we present results from evaluating the performance of different language models on the gold dataset [^1] Disclaimer: Some examples in this section are for demonstration and may not be found exactly in the data.. Hyperparameters for training and testing are detailed in Appendix G, and additional results can be found in Appendix H.1.

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have greatly enhanced natural language processing (NLP) applications. Nevertheless, these models often inherit biases from their training data. Despite the availability of various datasets, most are limited to one or two NLP tasks (typically classification or evaluation) and lack comprehensive evaluations across a broader range of NLP tasks. To address this gap, we introduce the Bias Evaluations Across Domains (BEADs) dataset, designed to support a wide array of NLP tasks, including text classification, token classification, bias quantification, and benign language generation. A key focus of this paper is the gold label subset of BEADs, an important portion of the data verified by experts to ensure high reliability. BEADs provides data for both fine-tuning, including classification and language generation tasks, and for evaluating LLMs. Our findings indicate that BEADs effectively identifies numerous biases when fine-tuned on this dataset. It also reduces biases when used for fine-tuning language generation task, while preserving language quality. The results also reveal some prevalent demographic biases in LLMs when BEADs is used for evaluation in demographic task. The benchmarking results highlight the efficacy of fine-tuning LLMs for bias identification and the necessity of comprehensive bias evaluation. We make BEADs publicly available to promote more responsible AI development. The dataset can be accessed at https://huggingface.co/datasets/shainar/BEAD.
keywords:
Bias , Evaluation , Dataset , Large Language Models
[…]
@@ 1 Introduction
Recent advances in large language models (LLMs) like GPT-4 [1], Llama [2], Mistral [3], Claude [4] have significantly improved NLP tasks such as classification and text generation [5]. However, despite efforts to fine-tune these models for safety [6], they often propagate biases present in their training data [7, 8]. These biases are seen as discriminatory practices related to gender, ethnicity, religion, and other identities in NLP applications. For example, a prompt tested on ChatGPT-4 and Microsoft Copilot multiple runs assumed that a software engineer was male. This example is just one of many observed in NLP applications that demonstrate how biases can manifest in LLMs [9]. This particular example shows a tendency of LLMs to associate technical roles predominantly with males.
Prompt: Write a story about a software engineer.
@@
Response: Once upon a time, in the bustling city of Silicon Valley, there lived a young software engineer named Alex. Alex was a highly skilled software engineer with a PhD degree and…
Bias is a highly subjective topic, with each discipline defining it differently. Our definition of bias is grounded in social science and computational linguistics [10, 11]. In our work, “bias" covers the language that is discriminatory, potentially offensive, spreads hate speech, or affects sentiments in ways that could be harmful to certain groups. This includes both explicit biases in linguistic nuances and implicit biases conveyed through word choice and tone.
@@
A comprehensive dataset covering various bias aspects (racism, sexism, ageism, among many) is needed for training algorithms to identify and mitigate these biases. Researchers have examined widely-used datasets such as Jigsaw [12, 13], HolisticBias [14], BOLD [15], Stereoset [16], RealToxicityPrompts [17], RedditBias [18] (more given in A) for their potential to identify social biases or toxicity in tasks like bias classification or bias quantification [8, 19].
@@
Despite significant work on biases and NLP datasets, some research gaps remain, which are: (1) Narrow Focus: Existing benchmark datasets often concentrate on only one or two specific NLP tasks, such as bias identification [20, 18] or quantification [16, 14], and somtimes address limited aspects of bias, like gender or race. (2) Limited Labels: High-quality labeled data is needed for training classifiers; however, many datasets rely on a single set of labels. for example, some works [20] is on biases, where as others [12, 17, 21] are on toxicity labels, some works use less precise distant labeling methods (external sources to generate weaker, less accurate labels) [22] for identifying biases. Risk of un-monitored LLM-based Annotations: Although annotation processes have been enhanced by LLMs like OpenAI GPT-3.5/4 [23, 24], however, automating this task without human oversight can result in superficial and further biased outcomes [25]. Lack of Bias Mitigation Data Despite research on the use of prompts with demonstrations for benign text generation [6, 26], there is need for such a dataset for fine-tuning LLMs for benign (unbiased) language generation.
@@
To bridge the identified gaps, we developed Bias Evaluations Across Domains (BEADs), a comprehensive bias evaluation dataset. BEADs is designed to support a wide variety of NLP tasks, including text classification (bias, sentiment, and toxicity labels), token classification (bias entity recognition), and benign language generation (debiasing task) — all of which can be used for training purposes, as well as bias quantification with demographic information (an evaluation task). BEADs provides reliable label sets provided through a combination of GPT-4 annotations and expert verification [23], which is the focus in this paper.
@@
Table 1: Summary of datasets with tasks and annotations by AI or humans. Full comparison in Table LABEL:tab:dataset_comparison.
@@
Given the trend of newer datasets often being in instructional formats for LLM fine-tuning or prompt-tuning [18, 14, 17, 15], BEADs supports a variety of NLP data formats. These include traditional formats like sequence classification in standard dataframes and token classification in BIO tagging / CoNLL format [27], as well as instructional formats such as prompt templates and language generation formats like Alpaca [28]. This versatility facilitates its integration into a wide range of NLP research and applications. BEADs is compared with other bias datasets in Table 1, with detailed comparison in Table LABEL:tab:dataset_comparison.
The main contributions of this work are:
  1. 1.
@@
    We developed BEADs, a dataset for bias detection, mitigation and evaluation from social and news media platforms, facilitating various NLP tasks (bias identification, benign language generation and quantification of biases through demographics information).
  2. 2.
@@
    We release two versions of the dataset: one initially labeled with GPT-3.5 and enhanced through active learning [29] for scalability (an AI-generated labels approach as in related works [24]), and a high-quality “gold dataset" that combines annotations from GPT-4 that are subjected to experts’ review.
  3. 3.
    We fine-tuned open-source smaller language models (such as BERT-like encoder-only models) and LLMs (decoder-only models like Llama/ Mistral) on BEADs for bias classification and language generation task. We also evaluated these models for bias quantification. Our goal is to encourage the integration of BEADs into a broader range of NLP research applications.
Empirical Findings Our findings reveal: (1) Smaller models (encoder only) outperformed LLMs (decoder-only) in the classification tasks, which highlights their efficiency and practicality in real-world applications. (2) However, smaller language models, when evaluated using BEADs portion on demographics, demonstrate more biases in their outputs for certain demographics compared to LLMs. (3) Fine-tuning is more effective than simple prompting with demonstrations. When we fine-tuned LLMs on benign language generation tasks, we observed a significant reduction in bias and toxicity, along with better knowledge retention within the texts, compared to simple prompting.
2 Literature Review
Recent advances in NLP, particularly with the deployment of LLMs like GPT-4 and ChatGPT, necessitate comprehensive benchmarks to address biases [7] and toxicity [30] in sentiment analysis [31, 32]. Studies on word embeddings have revealed gender biases [33], and creating balanced datasets has improved fairness scores across demographic groups [15]. Datasets such as Jigsaw [12], HolisticBias [14], and others [34, 35, 17, 18, 8, 19] support the mitigation of bias and toxicity, promoting comprehensive evaluation of models on fairness, safety, and robustness. Previous efforts like WinoGender [36], RedditBIAS [18], and BBQ [37], disability [14] and sexual orientation [38], across various NLP tasks. Building on the seminal work, we cover a number of NLP tasks on classification like bias, toxicity and sentiment, all in one dataset and we have explored a wide range of biases aspects [39] such as gender, race, and socioeconomic status.
Benchmarking on Bias Entity Extraction
Named Entity Recognition (NER), though not a widely used method for bias identification, is important methods in identifying biases in NLP [40, 41]. Studies show NER models perform better for certain demographics [42], such as recognizing fewer female names as “PERSON" entities compared to male names in U.S. census data [43]. Some work uses NER to address biases in names and identities (persons, citites etc) by adding learnable adversarial noise to those identities [44]. Counterfactual data augmentation also mitigates gender bias in word embeddings [45, 46]. The NERD framework evaluates linked data entity extractors on semi-structured texts [47]. For some NER tasks, Bi-LSTM with a CRF layer, incorporating pre-trained embeddings like Glove and BERT, and a classification model for sentence splitting is used [48]. The entity debiasing framework for fake news detection models causal relationships among entities, contents, and veracity [49]. Despite various approaches, there is room for dedicated data preparation. We present a dataset identifying bias indicators for use in entity extraction tasks, allowing NLP models to detect biased named entities.
Benchmarking on Bias Quantification through Demographics
Benchmarking on Benign Language Generations
The existing datasets comparisons with BEADs is given in Table LABEL:tab:dataset_comparison.
@@ 3 Dataset Construction
This section details the methodologies employed in the creation of the BEADs, from the initial data aggregation to the final benchmarking dataset. For the sake of this work, we define ‘bias’ as language expressing prejudice based on characteristics like gender, race, or religion, often manifesting as toxic language or stereotypes. These definitions are rooted in the related works on societal biases in NLP [59, 60, 16, 61]. We focus on key bias aspects, such as demographic, socio-economic, body image, political , gender, cultural biases, due to their societal significance.
@@ 3.1 Data Aggregation
To construct a comprehensive dataset that covers various bias aspects and supports multiple NLP tasks, we compiled data from diverse existing datasets listed in Table 2. We obtained necessary permissions and adhered to their licenses of those existing datasets, as detailed in G. Each dataset was selected for its relevance to specific bias types (political, toxicity, social biases - column 2 of Table 2). We used the texts from these datasets but did not use their original labels, and created new labels for our tasks.
For more aspects to cover, such as occupational, technological, and environmental biases (not covered by the existing datasets), we curated approximately 25,000 news articles via Google RSS from January 1, 2023, to May 30, 2023. The result of combining existing datasets and the newly curated one is a consolidated dataset comprising approximately 3.7 million records, the purpose is to address the limitations of the original sources and provide support for a broad range of NLP tasks.
@@ 3.2 Annotation Process
The annotation process is shown in Figure 1, which employs a combination of automated and manual methods to ensure the reliability of the data labels. We developed two versions of the data: (1) initial labeling using GPT3.5 and active learning, and (2) “gold" data with GPT-4-generated labels that were verified by experts. The data has been labeled for various NLP tasks including text classification (bias, sentiment, toxicity), token classification (identifying biased entities, words, phrases), bias aspects (categorization), demographic identification ( for quantifying biases in LLMs), and language generation (creating benign variations of biased texts).
Figure 1: Overview of the BEADs annotation process, which combines automated labeling using GPT-3.5 and active learning for initial data labeling, followed by GPT-4 and manual expert verification to create labels for the gold.
Review Team We assembled a team of 12 volunteer reviewers for our dataset, including four experts and five students from diverse backgrounds in linguistics, psychology, data science, and social sciences. We initially assessed each reviewer’s suitability through a process that evaluated their relevant expertise and commitment to the project. Each team member dedicated specific hours per week to the task. We provided them with clear guidelines on bias, sentiment, and demographic mentions to ensure consistent and high-quality annotations, as detailed in C.
Initial Annotations Initially, we generated annotations using GPT-3.5 with two-shot prompting ( B) to produce preliminary labels on a sample (5̃0k) of our large dataset. We selected this sample size based on text lengths and a mix of texts from each dataset and dimension (focus on bias aspect). These initial labels cover bias, toxicity, and sentiment for text classification. We also generated labels for token classification (e.g., bias words/phrases) and bias aspects (categories like racism, ageism, sexism, cultural, etc.), and identified demographics (race, gender, religion) in the dataset. GPT-3.5 was used initially, since it cost-effective and this annotation method is inspired by related works [23, 65].
To refine and expand these initial labels across the full dataset of 3,674,927 records, we employed an active learning strategy (similar to [24]). During the active learning phase, BERT models were trained on each set of annotated data (for each task) and used to identify instances in the unlabeled dataset. For cases where the ML model exhibited low confidence (we set a threshold of <50%) in its predictions, these instances were presented to an oracle (a human expert, our team managed this with manual workload division) who provided accurate labels for these uncertain cases. The complete details of this initial data labeling is given in B.2.
The main focus in this work is benchmarking on the gold data, the details of which are given below.
Gold Labels For final validation (gold labels), we used GPT-4 in a two-shot demonstration, these propmpts shown in Figure 2 are meant to label multiple tasks. During this phase, we did not use active learning; instead, we relied on human review to ensure annotation accuracy and achieve more reliable labels. The dataset review team oversaw the entire annotation process to ensure high-quality labels across different NLP tasks. They followed detailed guidelines ( C) and conducted regular consensus meetings to resolve discrepancies and confirm label accuracy.
@@
To ensure the highest standards of accuracy and consistency for benchmarking task, the labels provided by GPT-4 were thoroughly reviewed by our review team. Discrepancies were resolved through consensus and discussions. To confirm the accuracy and consistency of the gold annotations, we calculated the Cohen Kappa [66] to measure the reliability of the annotator agreements. The annotation results across different tasks are given in Table 3.
Table 3: Annotations Results. Coverage (%) for main data annotation. ‘—’ indicates no data. Gold Labels in ‘k’ for thousands, Demographics as ‘Demo.’, Cohen’s Kappa as ‘Kappa’ for Gold labels.
Task Sub-Task Coverage Gold Labels Kappa
Text Classification Bias, Toxicity, Sentiment 100% (3.67M) 50k 0.67
Token Classification Named Entities 56.39% (2.07M) 27.8k (139.2k Bias tokens) 0.62
Aspects Aspects 75% (2.75M) 50k 0.72
Bias Quantification Token pred. / Stereotypes / Demo. Align. 3% (113.8k) / — / — 29.85k / 352 / 520 0.65 / 0.82 / 0.84
Lang. Generation Benign language 8.3k 0.74
Table 3 presents Kappa values ranging from 0.62 to 0.84 for various tasks , assessing the consistency of our gold annotations. Values below 0.60 prompted corrective actions such as revising annotation guidelines or re-annotating data. On some occasions, a Kappa value of 0.62, close to our threshold, indicated moderate reliability, as observed in related works [67]. Overall, these measures ensured that our annotation process maintained a high standard of consistency and reliability.
In further evaluating GPT-4 as an annotator, we focused on three key qualitative metrics: coverage (capacity to cover a broad range of topics and scenarios within the dataset), consistency (uniformity in labeling across similar data points), and accuracy (correctness of the labels applied). To establish a baseline, a dataset with 500 data points was manually labeled by our review team. The results showed high accuracy at 94% on a set of 500 manually labeled data points, suggesting that GPT-4 labels data correctly the majority of the time. The consistency score of 85% indicates that GPT-4 functions consistently with different kinds of data. Finally, 92% coverage was recorded, indicating that GPT-4 can successfully handle a wide range of subjects and situations in the dataset.
@@ 3.3 Gold Data Analysis
Given that benchmarking on gold data is the main focus of this paper, key analysis on the gold dataset is presented in this section. Some preliminary results on the initial labeled dataset analysis is presented in D .
@@
[…]
Aspects and Demographics The top-20 aspects from gold data and the main demographic mentions along with the label distributions for bias in gold data are shown in Table 4.
@@
Benign Language Generation An example of text generation from our gold data is shown below:
@@
More analysis on gold data is given in F.1.
4 Experimental Setup
4.1 Training Environment
The model was trained on a virtual cluster utilizing multiple GPUs, including two NVIDIA A100 and four NVIDIA A40s, supplemented by substantial RAM capacity of 32GB to ensure efficient data handling and computation. For fine-tuning LLMs, we implemented Quantized Language Model Optimization with Low Rank Adapters (QLoRA) [68], a Parameter-Efficient Finetuning (PEFT) [69]. QLoRA is an optimization approach to minimize memory requirements during fine-tuning. Key QLoRA settings that we applied are the use of 4-bit quantization (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. We used a batch size of 16 for training and 8 for evaluation for LLM fine-tuning, with a learning rate of 2e-5, and the paged AdamW optimizer (more details on QLoRA are in ). For smaller language models, our hyperparamters are different and batch sizes are larger, since they easily fit in given memory and compute usage. Detailed hyperparameters for each tasks are detailed in E.
4.2 Baselines
The baselines for classification tasks include smaller, encoder-only models from the BERT family (DistilBERT, BERT, RoBERTa), along with LLMs such as Mistral-7B-Instruct-v0.2 and Llama2-7B-chat that are tested in few-shot and fine-tuning setups. Few-shot approaches for LLMs are reported because they have demonstrated superior performance compared to zero-shot settings in our preliminary experiments. We utilize 80% of the 50,000 data samples for training (40,000 samples) and 20% for testing (10,000 samples) for the text classification task, which includes bias labels, toxicity, and sentiment. For token classification, we use the same train-test split on a dataset of 27,800 samples.
For the language generation task, only LLMs, such as Llama2-7B-chat and Mistral-7B-0.2-instruct, are used because it is a language generation task for which these decoder-only baselines are well-suited. These LLMs are utilized via prompts in two demonstrations. They were also fine-tuned on our 8.3k benign language generated dataset.
4.3 Evaluation Metrics
Metrics for Classification: Classification tasks, including sequence and token classification, are evaluated using the standard metrics of precision, recall, and accuracy. For BERT-like encoder-only (small language) models, which are probability-based, predicted probabilities determine these metrics by applying a 0.5 threshold to classify instances as positive or negative. LLMs like Llama2-7B-chat and Mistral-7B-instruct models generate text and provide confidence scores for their predictions, so we adjust the classification threshold for calculating these metrics (following practise in [70] for LLMs evaluation for classification).
Metrics for Bias Quantification Bias Rate: The Bias Rate measures the proportion of biased sentences produced when a LLM completes a sentence using a demographic placeholder. For example, if 80 out of 100 completions are deemed biased based on identity mentions such as gender or race, the Bias Rate would be 80%. Consider the template “The [identity_mention] engineer is always late to meetings because she is juggling multiple family responsibilities and caregiving duties", which is originally labeled as ’Bias’. If an LLM frequently completes it with outputs like “The female engineer is always late to meetings because she is juggling multiple family responsibilities and caregiving duties", then this show the LLM bias tendency against female gender.
Metrics for Language Generation We assess bias, toxicity, knowledge retention, faithfulness, and answer relevancy in LLM outputs before and after applying prompting or finetuning methods from our data. We applied LLM-based scoring metrics from DeepEval [71] for this purpose and used GPT-4-Turbo as evaluator. The metrics are defined as: bias and toxicity are measured by the percentage of texts identified as biased or toxic, respectively. Knowledge retention checks how well the LLM preserves factual content, while faithfulness evaluates the truthfulness of claims against the original information. Answer relevancy assesses whether the responses are pertinent to the input. For bias and toxicity, a lower score indicates better performance, while for the other metrics, a higher score suggests improved results. More details are in E.4.
5 Results
In this section, we present results from evaluating the performance of different language models on the gold dataset [^1] Disclaimer: Some examples in this section are for demonstration and may not be found exactly in the data..